OpenStudy (anonymous):

what is the domain of the function the square root of 64-x^2

OpenStudy (anonymous):

Quantities under root sign can't be less than 0. Set >= 0: \[64-x^2 \ge 0 \] \[64 ≥ x^2\] \[64 ≥ x^2\] \[x \le 8\]

OpenStudy (anonymous):

sorry, there is a bit more. Because you end up with a + and - when you solve a squareroot you should instead get for the domain: . . \[64 \ge x^2\] \[x \le \left| 8 \right| \] \[-8 \le x \le 8\]